OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 4x + 7

OpenStudy (anonymous):

y = (x + 2)2 + 3 i believe this is the answer, i just want conformation

OpenStudy (jdoe0001):

\(\bf y=ax^2+bx+c \textit{ ,standard form}\\ y=a(x-h)^2+k \textit{ ,vertex form}\)

OpenStudy (jdoe0001):

so you'd need to "complete the square" for that

OpenStudy (anonymous):

so would i be correct?